OpenStudy (surana):

The graph of y>x^2 -6x+7 contains which of the following points?

OpenStudy (surana):

^2 represents exponent of 2. Answers are... A: 0,0 B: -3, 2 C: 5,-1 D: 3,-1

OpenStudy (anonymous):

Have you graphed it?

OpenStudy (anonymous):

For question !, graph it like i had you do the last one =)

OpenStudy (surana):

Trying to right now.

OpenStudy (anonymous):

D is your answer =)

OpenStudy (surana):

Could you explain please?

OpenStudy (anonymous):

Yes, if you graph the inequality, that is the only point from your answers that is inside the shaded region

OpenStudy (campbell_st):

the question is looking for more than 1 point... a simple process is to use substitution using the point (0, 0) or x = 0, y = 0 substitute them and \[0 > 0^2 - 6 \times 0 + 7\] is this true..? if it is, then its one the the correct answers

OpenStudy (anonymous):

Guess there is no need. These have to be the vertices. D would be the answer. one of the ways of find x and y - vertices is : where b is the value attached to the x variable, a is attached to the x^2 value and c is the one with no variable. The x value will be 6/2=3 and y will be (6^2 -4*7)/4=2